OpenStudy (anonymous):

If ƒ(x ) = x 2 + 1 and g(x ) = 3x + 1, find ƒ(-1) + 5g(-3). -40 -38 7

OpenStudy (anonymous):

@imqwerty

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Have you considered substituting the given values and evaluating the expression?

OpenStudy (anonymous):

well i mean im in online classes and no teacher to really give me a better concept of the work

OpenStudy (tkhunny):

Example h(x) = 3x -1 h(1) = 3(1) - 1 = 3-2 = 2 h(-2) = 3(-2) - 1 = -6-1 = -7 h(7) = 3(7) - 1 = 21 - 1 = 20 That's all you need. That's what function notation does.